OpenStudy (anonymous):

4. Use the function rule ƒ(x) = x • |–1|x. Find the output ƒ(2). (1 point) 4 3 –2 2

OpenStudy (solomonzelman):

plug in 2 for x.

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm ƒ(2) = 2 \times |–1|2=2\times (1)2=2\times 2 =4 }\)